OpenStudy (anonymous):

order the following fractions from smallest to largerst: 3/8 , 1/3 , 5/9 , 2/5 .

OpenStudy (tkhunny):

One COULD make them all the same denominator, but that would require finding the LCM for 8, 3, 9, and 5. That's not TERRIBLE, but another way presents itself on this one. Just create short decimal approximations for each. 3/8 = 0.????? 1/3 = 0.?????... 5/9 = 0.?????... 2/5 = 0.?????

OpenStudy (anonymous):

so the order is 1/3 , 3/8 , 2/5 , 5/9

OpenStudy (tkhunny):

I do wish you would provide intermediate results. That's the only way to know you didn't just copy the answer out of a book.